Thanks again!!!!
Solve by addition. If a unique answer does not exist, state whether it is inconsistent or dependant. McGraw Hill. 6th editon. 8.2. #20.
x+5y=10
-2x-10y=20

Answer by stanbon(48516)

1st: x+5y=10
2nd: -2x-10y=20
Multiply 1st by -2 to get:
3rd: -2x-10y=-40
Notice that 3rd contradict 2nd.
If you subtract 3rd from 2nd you get:
0=60
That shows the original system is a
contradiction. That means there is
no solution. It also means the graphs
of the two lines are parallel.
cheers,
Stan H.